6x^(2)+19x-36. Would you show me how to do it please This is algebra factoring
Answer by ExplanationCentral(10)

The expression is in quadratic form.
The product of the roots must be -36/6, and the sum of the roots should be -19/6.
So by solving a + b = -19/6 and a*b = -36/6, we get the two roots.
a = -9/2 and b = 4/3
The quadratic would be (x+(9/2))(x-(4/3)).